OpenStudy (shaleiah):

A salsa recipe uses green pepper, onion, and tomato in the extended ratio 2 : 5 : 8. How many cups of onion are needed to make 105 cups of salsa?

OpenStudy (shaleiah):

@Michele_Laino

jhonyy9 (jhonyy9):

green pepper onion tomato 2 5 8 for one cup x for 105 cups x = ?

jhonyy9 (jhonyy9):

5 ---- 1 cup x ---- 105 cups ------------------ x = 105*5/1 = ?

OpenStudy (shaleiah):

525

jhonyy9 (jhonyy9):

us the rule of 3 simple

OpenStudy (michele_laino):

I write the proportion, in this way: \[\Large p:o:t = 2:5:8\] so, if I apply a property of proportions, I get: \[\Large \left( {p + o + t} \right):o = \left( {2 + 5 + 8} \right):5\] or this one: \[\Large 105:o = 15:5\] being \(o+p+t=105\) please solve for \(o\)

OpenStudy (anonymous):

105/5=21 21 cups of onions

OpenStudy (michele_laino):

If I apply the fundamental property of proportions, I get: \[o = \frac{{105 \times 5}}{{15}} = ...?\]

OpenStudy (michele_laino):

hint: \[o = \frac{{105 \times 5}}{{15}} = \frac{{525}}{{15}} = ...?\]

OpenStudy (shaleiah):

35 cups

OpenStudy (michele_laino):

please wait a moment I'm checking my computation...

OpenStudy (michele_laino):

I'm very sorry, I have made a wrong reasoning

OpenStudy (shaleiah):

:(

OpenStudy (michele_laino):

please wait a moment, I'm thinking...

jhonyy9 (jhonyy9):

@Michele_Laino what is your opinion about above wrote my idea using the rule of 3 simple 1 cup salsa 5 onion 105 cup salsa x onion ------------------- x = 105*5/1 = ?

jhonyy9 (jhonyy9):

i think this is the best easy way

OpenStudy (michele_laino):

we can write this expression: \[2N + 5N + 8N = 105\] or: \[15N = 105\] where \(N\) is a quantity to be determined please solve for \(N\)

OpenStudy (shaleiah):

7

OpenStudy (michele_laino):

ok! So, the requested cups of onion are: \(5 \times 7=...?\)

OpenStudy (shaleiah):

35

OpenStudy (michele_laino):

I don't know, since I'm not sure it is a 3 simple problem @jhonyy9 that's right! @shaleiah

OpenStudy (michele_laino):

more precisely we have \(2 \times 7=14\) cups of green pepper and \(8 \times 7=56\) cups of tomato

jhonyy9 (jhonyy9):

@Michele_Laino sorry but i think now you are not right sure than use 2 5 8 ratio for one cup salsa 2 cup 5 cup => these for one cup salsa 8 cup so than for 105 cup salsa you need using 5*105 cup onion sure

OpenStudy (michele_laino):

here is my reasoning: 1 cup of salsa contains \(2/ 15\) of cup of peppers, \(5/15\) of cup of onions, and \(8/15\) of cup of tomato so 105 cups of salsa contains \(105 \times (5/15)=35\) cups of onions

OpenStudy (michele_laino):

@jhonyy9

jhonyy9 (jhonyy9):

yes @Michele_Laino you are right sure with this reason - now i see it - this way of 3 simple not is the right way here - thank you for your help - great -

OpenStudy (michele_laino):

thanks! :) @jhonyy9